A man age is now his son’s age in reverse. His son was born when he was 45. How old is his son?

Answer by ankor@dixie-net.com(22740)

A man age is now his son’s age in reverse.
His son was born when he was 45. How old is his son?
:
Let x = the 10's digit
Let y = the units digit
:
10x + y = 10y + x + 45
10x - x = 10y - y + 45
9x = 9y + 45
Simplify divide by 9
x = y + 5
Several options here, we know y < 4, and x > 5
Man is 72, son is 27
or
Man is 61, son is 16
or
Man is 83, son is 38
or
Man is 94, son is 49